### v1

# TelePhantom: A User-Friendly Teleoperation System with Virtual Assistance for Enhanced Effectiveness

Abstract
Dexterous manipulation is a critical area of robotics. In this field, teleoperation faces three key challenges: user-friendliness for novices, safety assurance, and transferability across different platforms. While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation. To address these limitations, we introduce TelePhantom. This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000. TelePhantom allows the user to see a virtual robot that represents the outcome of the user’s next movement. By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely. We demonstrate its superiority over other teleoperation systems across five tasks, emphasize its ease of use, and highlight its ease of deployment across diverse input sensors and robotic platforms. We will release our code and a deployment document on our website https://telephantom.github.io/.
{strip}
Figure 1: TelePhantom is a user-friendly teleoperation system enabling the real-time virtual preview before robot execution.
I Introduction
Figure 2: Overview of the TelePhantom System Architecture: Our system consists of three main components: (1) Various input devices including RGBD cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a unified interface to process wrist and hand poses data, feeding into our Phantom System with joint-to-joint mapping and non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
Teleoperation has long been recognized as an essential component in gathering on-robot data for learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data. However, effectively controlling robots with dexterous hands remains a significant challenge. The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex. These inherent structural and functional differences, combined with the precision required for precise manipulation tasks, often result in inefficient and potentially compromised data collection processes.
Existing teleoperation systems face three significant challenges. First, novice users often struggle to control the robot effectively due to the lack of intuitive feedback on how their commands translate to robot actions, especially during complex manipulation tasks. Second, without proper safeguards and real-time guidance, users may inadvertently issue unsafe commands that could damage the robot or its surroundings. Third, most current systems are tightly coupled with specific input devices or robot platforms, limiting their adaptability and broader application. These challenges significantly impact the quality and efficiency of data collection for robot learning.
To address these challenges, we introduce TelePhantom, a teleoperation system that delivers precise control while maintaining affordability through its sub-\$1,000 hardware implementation. At its core, TelePhantom provides an interface where a virtual robotic arm spatially overlaps with the physical arm, allowing users to preview and confirm intended movements through an external input device (e.g., a foot pedal) before execution on the physical system.
Building upon this core design, our modular framework (Figure 2) addresses key challenges in human-robot motion mapping through the seamless integration of multiple components. Our system architecture utilizes IMU-based motion capture systems and mocap gloves to estimate wrist (Section III-A) and hand poses (Section III-B1). We use the SMPL-X standard (Section III-A1) to bridge all modules together, enabling flexible integration while maintaining system cohesion. These inputs are mapped to the robotic hand using our joint-to-joint retargeting approach (Section III-B2) and non-collision retargeting module (Section III-C). We design each module to be independent yet interoperable, enabling flexible system configuration while maintaining robust performance.
We present the major contributions of TelePhantom to advance the field of teleoperation:
  • Low-Cost System: We introduce an economical yet high-performance teleoperation system that integrates inertial motion capture technology and mocap gloves, with a total hardware cost under \$1,000, making it affordable for research.
  • Interactive Visual Assistance: We develop a visualization interface that enables users to preview intended actions before execution, enhancing teleoperation precision and accessibility for both novice and experienced users during complex tasks.
  • Versatile Modules: We design a modular assisted teleoperation pipeline that bridges various input devices with different robot end-effectors through joint-to-joint mapping and non-collision retargeting. This cross-platform design enables the versatile module to be integrated with diverse input methods and robots, making it a flexible solution for various manipulation scenarios.
II Related Work
II-A Robot Teleoperation Frameworks for Manipulation
The growing interest in training robots via imitation learning has driven the need for extensive, high-quality real robot datasets. Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks. Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [12, 13, 14], wearable gloves [15, 11, 16, 17], exoskeletons [18, 19], and other low-cost devices, each offering distinct benefits in terms of accessibility, precision, and generalizability.
While vision-based teleoperation methods offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands. Although specialized hardware [4] can improve accuracy, it substantially increases system costs. Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs. Many systems [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control. Furthermore, their customization options typically require time-consuming manual parameter tuning that demands significant expertise.
Figure 3: Pipeline of Our Assisted Module: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution. We achieve precise alignment between virtual and physical robots through AprilTag calibration.
II-B Visual Feedback in Teleoperation Systems
Visual feedback plays a crucial role in teleoperation systems by providing users with spatial awareness and task-relevant information. Traditional approaches primarily rely on real-time camera feeds [20, 21], which can be limited by occlusions and restricted viewing angles. Moreover, the lack of preview capabilities for intended commands leaves users (especially novice users) uncertain about the robot’s response, which can increase error rates and user fatigue.
To address these limitations, researchers have explored enhanced visualization techniques [22, 23, 24] that overlay virtual information onto the workspace view. While these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily focus on visualizing task-specific elements. This focus on task-specific visualization leaves users uncertain about how the robot will execute its commands.
III Modular Teleoperation Framework
Our teleoperation system consists of three main components: wrist pose estimation via IMU-based tracking, hand pose estimation using a motion capture glove, and non-collision retargeting for safe robot command generation. The following subsections detail each component of this pipeline.
III-A Wrist Pose Estimation
III-A1 SMPL-X Standard
We adopt the SMPL-X standard [25] as our unified human representation. The SMPL-X model represents a kinematic tree with standardized joint coordinate systems, where each joint’s pose describes a rotation and translation relative to its parent frame. This hierarchical structure enables consistent pose representation across different motion capture devices and simplifies the integration with our TelePhantom system.
III-A2 IMU-based Tracking System
While many researchers adopt vision-based motion capture systems in teleoperation, these systems require multiple cameras, precise calibration, and controlled lighting conditions, which limits their practical deployment. IMU-based motion capture offers robust tracking regardless of visual conditions and occlusions.
Our system uses IMU sensors located on the user’s palm, forearm, and upper arm. Given the three IMUs, we utilize the REBOCAP framework [26] to obtain the specific SMPL-X model of the human body [25, 27]. Following the SMPL-X convention, we define the world frame system at the midpoint between the feet. All positions and orientations in this paper are expressed in this coordinate system unless otherwise specified. After obtaining the SMPL-X model, we compute the wrist position $\mathbf{p}_{w}\in\mathbb{R}^{3}$ and orientation $\mathbf{R}_{w}\in SO(3)$ relative to this world frame.
Based on the wrist parameters, we map the user’s wrist pose to the end-effector pose through a transformation function $\mathcal{T}$. Specifically, we use the relative wrist position to control the end-effector position and the absolute wrist orientation to control the end-effector orientation:
$\mathbf{p}_{e}(t)=\mathbf{p}_{e}(0)+(\mathbf{p}_{w}(t)-\mathbf{p}_{w}(0)),\quad\mathbf{R}_{e}(t)=\mathbf{R}_{w}(t),$ (1)
where $\mathbf{p}_{e}(t)$ and $\mathbf{R}_{e}(t)$ represent the end-effector position and orientation at time $t$, and $\mathbf{p}_{w}(t)$ and $\mathbf{R}_{w}(t)$ represent the user’s wrist position and orientation at time $t$.
III-B Hand Pose Estimation
III-B1 Human Hand Joint Value Acquisition
Vision-based methods typically use deep learning models to predict 3D joint positions $\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ from RGB or RGB-D images. Due to the projective nature of imaging, vision-based systems struggle with self-occlusion, forcing researchers to infer joint positions from priors rather than direct observation. In contrast, mocap gloves [28] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors despite their higher hardware cost. For our teleoperation scenario, we opt for the mocap-based approach to ensure robust hand tracking.
III-B2 Joint-to-Joint Mapping
We propose a direct joint-to-joint mapping function $\mathcal{M}:\mathbb{R}^{27}\rightarrow\mathbb{R}^{16}$ that transforms mocap glove readings to LEAP hand [29] joint configurations. Let $q_{g}\in\mathbb{R}^{27}$ denote the glove joint values and $q_{r}\in\mathbb{R}^{16}$ denote the robot joint values. For each robot joint $i$, we define a mapping:
$q_{r}^{i}=f_{i}(q_{g}^{k_{i}}),\quad i\in{1,\ldots,16},$ (2)
where $k_{i}$ is the corresponding glove joint index for robot joint $i$ which is manually picked, and $f_{i}$ is a linear transformation:
$f_{i}(x)=s_{i}(x-b_{i})r_{i}.$ (3)
Here, $s_{i}\in\mathbb{R}^{+}$ is a scaling factor that normalizes the joint ranges, $b_{i}\in\mathbb{R}$ is a bias term that aligns the neutral positions, and $r_{i}\in$-1,1} is a direction indicator that ensures consistent joint rotations. This mapping guarantees:
$\max q_{g}^{k^{i}}\Leftrightarrow\max q_{r}^{i},\quad\min q_{g}^{k^{i}}\Leftrightarrow\min q_{r}^{i},$ (4)
Where $q_{g}$ and $q_{r}$ represent the joint angles of the glove and the Leap Hand, respectively, and max and min denote their upper and lower joint limits.
III-C Non-Collision Retargeting Network
In robotic manipulation systems, ensuring non-collision during operation is crucial for both physical simulator accuracy and hardware safety. Self-collisions can lead to unstable behavior in simulators and, more critically, cause hardware damage in real-world deployments. This challenge is particularly significant for high-DoF systems like the LEAP hand [29], where collisions often occur due to differences in action spaces and imprecise retargeting. As shown in Fig.4, direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig.4 (a). While traditional collision avoidance methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
To address these challenges, we develop a learning-based framework with two key components: (1) a Self-Collision Prediction Network (CPN) that predicts collision probabilities for each joint configuration and (2) a Configuration Correction Network (CCN) that maps invalid configurations to their closest valid counterparts. The CCN is trained with a composite loss function that balances configuration similarity and collision avoidance, ensuring both motion continuity and physical feasibility. As demonstrated in Fig. 4 (b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz). We provide the detailed technical implementation of this Pos-to-Pos Non-Collision Module in the Appendix.
(a) Current retargeting method [6] causing collision.
(b) Our Self-collision Avoidance Retargeting.
Figure 4: Comparison of hand configuration retargeting methods: (a) shows the direct mapping between human and robot hands leading to self-collision; (b) demonstrates our collision-aware retargeting approach that maintains safe configurations.
IV Phantom System
The phantom system aims to provide safe and intuitive robot teleoperation through a phantom-based motion preview mechanism. The process consists of three key steps: (1) the user’s motion commands are first rendered on a phantom robot, (2) the phantom is precisely aligned with the real robot using AprilTag markers [30], and (3) upon user confirmation via a foot pedal, the validated motion is executed by the physical robot. This preview-then-execute workflow, as shown in Fig. 3, enables users to verify and adjust planned motions before actual execution, significantly improving teleoperation safety and efficiency.
IV-A Why Phantom System Assistance?
IV-A1 Current Challenges
Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning. During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories.
In our hanging insertion experiments, we observed that users using direct teleoperation required an average of 4 corrective movements per attempt. These adjustments typically involved probing movements to locate the socket edge and multiple approach angle refinements before successful insertion. Such trial-and-error behavior introduced significant noise into the collected datasets. This issue is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
IV-A2 Proposed Solution
Our phantom-assisted teleoperation system addresses these data quality challenges through a novel preview mechanism. By visualizing a virtual “phantom” representation of the robot’s intended motion before executing physical movements, users can preview and refine their commands before committing to actual robot motion. This preview capability eliminates the need for exploratory movements in the physical space, as the real robot remains stationary until the user confirms the phantom pose. This design ensures that recorded trajectories contain only intentional, task-relevant actions without the noise of exploratory movements, directly benefiting downstream imitation learning applications with higher-quality demonstration data.
IV-B Technical Implementation of Phantom System
IV-B1 Sim-to-Real Alignment Using AprilTag
To maintain precise alignment between the phantom and physical robot across multiple camera viewpoints, we implement a robust calibration system using AprilTags [30]. Our approach establishes reliable reference frames through hand-eye calibration, and AprilTag pose estimation, linking both fixed and floating camera coordinate systems to the robot’s base frame.
Figure 5: Our transformation relationship: The number in the circle denotes the order of transformation acquisition.
As shown in Fig. 5, our calibration procedure involves multiple sequential transformations. We begin with hand-eye calibration of the fixed top-view camera (1), followed by detecting the AprilTag’s pose from this camera’s perspective (2). Using these relationships, we can compute the transformation between the AprilTag and the robot base frame (2). For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3). This chain of transformations ensures consistent phantom visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
IV-B2 Robot Phantom Visualization
Building upon this spatial alignment system, we implement the visual component of our phantom system. As shown in Fig. 6(a), we use the pyrender [31] library to render the virtual robot through the previous float camera. The rendered phantom image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment. This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the phantom accurately reflects the robot’s intended configuration in the workspace. The resulting overlay provides users with an intuitive preview of planned motions.
IV-B3 Multi-view Phantom Visualization
Task TelePhantom Open Teach [3] AnyTeleop[6] Telekinesis[32]
Pick & Place 1.0 0.8 1.0 0.9
Hang 0.9 - - -
Pour 1.0 0.8 0.7 0.7
Box Rotation 1.0 - 0.6 0.6
Cup Stacking 1.0 - 0.7 0.3
TABLE I: Real Robot Teleoperation Results. For the baseline methods, we use success rate as a metric reported in their papers. Especially, we use the success rate of an expert as a metric for Open Teach [3].
(a) Real-time phantom rendering
(b) Third-person view
(c) Top-down view
Figure 6: Rendering and Multi-view visualization system.
Task Success Rate Average Success Execution Time (s)
w/o phantom w/ phantom Difference $\uparrow$ w/o phantom w/ phantom Difference $\downarrow$
Pick & Place 0.6 1.0 +0.4 23.56 13.55 -10.01
Hang 0.6 1.0 +0.4 29.30 30.83 +1.53
Pour 0.9 0.8 -0.1 43.20 36.13 -7.07
Box Rotation 0.6 0.8 +0.2 30.47 19.12 -11.35
Cup Stacking 0.5 1.0 +0.5 31.54 18.91 -12.63
TABLE II: Effect of Phantom Assistance on New User Performance.
Our phantom system supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks. As shown in Fig. 6(b) and Fig. 6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives. For instance, during manipulation tasks, users can simultaneously monitor the robot’s vertical position from the third-person view while checking its planar alignment from the top-down view.
IV-B4 IO Control for Phantom Visualization
Figure 7: State transition of phantom control
Our system uses a foot pedal as an intuitive control interface for state transitions. As shown in Fig. 7, when the IO is activated, the physical robot stops while the phantom appears and responds to user commands for motion preview. Upon IO deactivation, the phantom disappears, and the system extracts its final configuration as the target pose, bypassing any exploratory movements made during the preview. This target is then processed by the mplib motion planning library [33] to generate an optimal trajectory for execution. After reaching the target pose, the robot automatically resumes real-time teleoperation until the next IO is activated. This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
V Experiments
Our experiments aim to address the following questions:
(Q1) How effective is TelePhantom vs Baselines?
(Q2) How beneficial is TelePhantom for new users?
(Q3) How versatile is TelePhantom on robots?
V-A Experiment Setup
TelePhantom integrates several key components. The user’s movements are captured via an inertial motion capture suit and a pair of strain gauge-based mocap gloves, providing high-fidelity joint angle data for both body posture and delicate hand movements. To visualize robot action before execution, we employ a RealSense RGB-D camera and render the phantom either on a standard 2D display or within a VR headset (e.g., Meta Quest 3). For the physical hardware, we use a UFactory xArm-6 robotic arm outfitted with a Leap hand [29], utilizing AprilTag [30] to align the initial poses of the phantom and the real robot.
Following the methodology outlined in [32, 6], users attempted each task 10 times. We reference the baseline success rates from their papers.
V-B Real Robot Teleoperation
We test TelePhantom on five real-world tasks that highlight different manipulation challenges:
  • Pick & Place: Grasp an object and place it at precise locations.
  • Pour: Tilt and rotate a cup to pour the beans into another container.
  • Hang: Hang a spoon on a peg, requiring both fine positioning and subtle wrist rotations.
  • Box Rotation: Rotate a box to change its orientation.
  • Cup Stacking: Stack a cup onto another cup.
Performance Comparison with Baselines. Addressing Q1 (effectiveness vs. baselines), we compare the real robot teleoperation performance of TelePhantom, operated by an expert, with Telekinesis [32], AnyTeleop [6] and Open Teach [3]. As summarized in Table I, TelePhantom outperforms baseline methods in terms of success rate. The baseline methods are limited to simple, short-horizon tasks. While the baseline methods can handle basic tasks, they fail to accurately capture subtle wrist rotations and movements in more delicate operations like the hung task. The superior performance of TelePhantom in these scenarios demonstrates its capability to handle more challenging tasks effectively.
Effect of Phantom. To address Q2 (benefits for new users), we assess TelePhantom’s utility by measuring both the success rate and the average execution time on the real robot (demonstration time in imitation learning) with and without phantom. As shown in Table II, TelePhantom significantly improves the performance of new users by improving their task success rate. The execution time on a real robot corresponds to the demonstration time required for imitation learning. High-quality demonstrations are typically associated with shorter execution times. With the assistance of TelePhantom, new users achieve a lower average execution time, indicating that they can produce higher-quality demonstrations compared to users without access to TelePhantom. For new users, command visualization and reduced execution time not only decrease physical strain but also improve safety.
V-C Deployment on Various Robots
TelePhantom’s modular design enables seamless adaptation to different input devices and end-effectors. Thanks to our robot-centric phantom approach and standardized transformation framework, users can readily configure the system for various robotic platforms by simply updating the kinematic parameters and transformation chain. This adaptability across different robot configurations, as posed in Q3, demonstrates TelePhantom’s versatility as a flexible teleoperation solution that can serve diverse robotic applications with minimal modification requirements.
VI Conclusion
In this paper, we presented TelePhantom, a low-cost teleoperation system featuring AR phantom feedback designed for deployment across diverse robotic platforms. Our experiments demonstrate that TelePhantom offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks. Through its innovative phantom visualization approach, the system provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the phantom and scene objects. Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.
Addressing these visualization challenges could yield several essential benefits. Improved occlusion handling would reduce user uncertainty during complex manipulation tasks, leading to higher-quality demonstration data for imitation learning. Additionally, more accurate spatial rendering would create a more intuitive teleoperation experience, potentially reducing user training time and cognitive load. These improvements would further advance TelePhantom’s goal of providing accessible, effective teleoperation capabilities across diverse robotic applications.

### v4

@@
-TelePhantom: A User-Friendly Teleoperation System with Virtual Assistance for Enhanced Effectiveness
+TelePreview: A User-Friendly Teleoperation System with Virtual Arm Assistance for Enhanced Effectiveness
 Abstract
-Dexterous manipulation is a critical area of robotics.
-In this field, teleoperation faces three key challenges: user-friendliness for novices, safety assurance, and transferability across different platforms.
+Teleoperation provides an effective way to collect robot data, which is crucial for learning from demonstrations.
+In this field, teleoperation faces several key challenges: user-friendliness for new users, safety assurance, and transferability across different platforms.
 While collecting real robot dexterous manipulation data by teleoperation to train robots has shown impressive results on diverse tasks, due to the morphological differences between human and robot hands, it is not only hard for new users to understand the action mapping but also raises potential safety concerns during operation.
-To address these limitations, we introduce TelePhantom.
+To address these limitations, we introduce TelePreview.
 This teleoperation system offers real-time visual feedback on robot actions based on human user inputs, with a total hardware cost of less than \$1,000.
-TelePhantom allows the user to see a virtual robot that represents the outcome of the user’s next movement.
+TelePreview allows the user to see a virtual robot that represents the outcome of the user’s next movement.
 By enabling flexible switching between command visualization and actual execution, this system helps new users learn how to demonstrate quickly and safely.
-We demonstrate its superiority over other teleoperation systems across five tasks, emphasize its ease of use, and highlight its ease of deployment across diverse input sensors and robotic platforms.
-We will release our code and a deployment document on our website https://telephantom.github.io/.
+We demonstrate that it outperforms other teleoperation systems across five tasks, emphasize its ease of use, and highlight its straightforward deployment across diverse robotic platforms.
+We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview/.
 {strip}
-Figure 1: TelePhantom is a user-friendly teleoperation system enabling the real-time virtual preview before robot execution.
+Figure 1: TelePreview is a user-friendly teleoperation system enabling the real-time virtual preview before robot execution.
 I Introduction
-Figure 2: Overview of the TelePhantom System Architecture: Our system consists of three main components: (1) Various input devices including RGBD cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a unified interface to process wrist and hand poses data, feeding into our Phantom System with joint-to-joint mapping and non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
-Teleoperation has long been recognized as an essential component in gathering on-robot data for learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
+Figure 2: Overview of the TelePreview System Architecture: Our system consists of three main components: (1) Various input devices, including RGB-D cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a standard to process wrist and hand pose data, feeding into our Preview System with joint-to-joint mapping and a non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
+Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
 An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data.
@@
 The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex.
-These inherent structural and functional differences, combined with the precision required for precise manipulation tasks, often result in inefficient and potentially compromised data collection processes.
+These inherent structural and functional disparities, combined with the precision required for fine-grained tasks, often result in inefficient and potentially compromised data collection processes.
 Existing teleoperation systems face three significant challenges.
-First, novice users often struggle to control the robot effectively due to the lack of intuitive feedback on how their commands translate to robot actions, especially during complex manipulation tasks.
+First, new users often struggle to control the robot effectively due to the lack of intuitive feedback on how their commands translate to robot actions, especially during complex manipulation tasks.
 Second, without proper safeguards and real-time guidance, users may inadvertently issue unsafe commands that could damage the robot or its surroundings.
-Third, most current systems are tightly coupled with specific input devices or robot platforms, limiting their adaptability and broader application.
+Third, most current systems are tightly coupled with specific input devices or robot platforms, limiting their adaptability.
 These challenges significantly impact the quality and efficiency of data collection for robot learning.
-To address these challenges, we introduce TelePhantom, a teleoperation system that delivers precise control while maintaining affordability through its sub-\$1,000 hardware implementation.
-At its core, TelePhantom provides an interface where a virtual robotic arm spatially overlaps with the physical arm, allowing users to preview and confirm intended movements through an external input device (e.g., a foot pedal) before execution on the physical system.
-Building upon this core design, our modular framework (Figure 2) addresses key challenges in human-robot motion mapping through the seamless integration of multiple components.
-Our system architecture utilizes IMU-based motion capture systems and mocap gloves to estimate wrist (Section III-A) and hand poses (Section III-B1).
-We use the SMPL-X standard (Section III-A1) to bridge all modules together, enabling flexible integration while maintaining system cohesion.
-These inputs are mapped to the robotic hand using our joint-to-joint retargeting approach (Section III-B2) and non-collision retargeting module (Section III-C).
-We design each module to be independent yet interoperable, enabling flexible system configuration while maintaining robust performance.
-We present the major contributions of TelePhantom to advance the field of teleoperation:
-  • Low-Cost System: We introduce an economical yet high-performance teleoperation system that integrates inertial motion capture technology and mocap gloves, with a total hardware cost under \$1,000, making it affordable for research.
-  • Interactive Visual Assistance: We develop a visualization interface that enables users to preview intended actions before execution, enhancing teleoperation precision and accessibility for both novice and experienced users during complex tasks.
-  • Versatile Modules: We design a modular assisted teleoperation pipeline that bridges various input devices with different robot end-effectors through joint-to-joint mapping and non-collision retargeting.
-This cross-platform design enables the versatile module to be integrated with diverse input methods and robots, making it a flexible solution for various manipulation scenarios.
+To address these issues, we introduce TelePreview, a teleoperation system that delivers precise control while maintaining affordability through its sub-\$1,000 hardware setup, as shown in Fig.
+2.
+At its core, TelePreview provides an interface that overlays a virtual robotic arm onto the real-world scene, enabling users to preview motions.
+Precisely, a foot pedal or other external signal IO allows for mode switching between preview (virtual-only) and align (physical action alignment) modes.
+In preview mode, users can verify and refine their intended actions with a virtual robot arm that is spatially aligned with the physical robot, ensuring that the final commands are safe and accurate once applied in the real world.
+We present the significant contributions of TelePreview to advance teleoperation:
+  • Interactive Visual Assistance: We develop a visualization interface that enables users to preview intended actions before execution, enhancing teleoperation precision and accessibility for both new and experienced users during complex tasks.
+  • Low-Cost System: We introduce an economical yet high-performance teleoperation solution that integrates inertial motion capture technology and mocap gloves at a total hardware cost of under \$1,000, making it affordable for research.
+  • Easy Adaptation to New Hardware: We design our approach so that migrating to different input or output devices requires adjusting only a small set of physically interpretable parameters, ensuring robust adaptation across diverse hardware setups.
 II Related Work
+Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution.
+We achieve precise alignment between virtual and physical robots through AprilTag calibration.
 II-A Robot Teleoperation Frameworks for Manipulation
-The growing interest in training robots via imitation learning has driven the need for extensive, high-quality real robot datasets.
-Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks.
-Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [12, 13, 14], wearable gloves [15, 11, 16, 17], exoskeletons [18, 19], and other low-cost devices, each offering distinct benefits in terms of accessibility, precision, and generalizability.
-While vision-based teleoperation methods offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands.
+The growing interest in training robots through imitation learning has driven the need for extensive, high-quality real robot datasets.
+Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks[13, 14, 15].
+Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [16, 17, 18], wearable gloves [19, 11, 20, 21], exoskeletons [22, 23], each offering distinct benefits in terms of accessibility, precision, and generalizability.
+Vision-based teleoperation systems [6] typically employ RGB or RGB-D cameras to detect hand poses.
+The hand pose retargeting module maps the human hand pose data obtained from perception algorithms into joint positions of the dexterous hand.
+We formulate this process as an optimization problem that minimizes the difference between the key point vectors of the human hand and the dexterous hand.
+To generate the end-effector motions, 3D positions of key points in the camera frame are calculated by using an RGB-D camera.
+The wrist pose is then computed by keypoint positions in the local wrist frame and global camera frame using the Perspective-n-Point (PnP) algorithm.
+Although vision-based teleoperation systems offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands.
+This limitation stems from the requirement of vision-based approaches for operators to keep their hands within the camera’s field of view, thereby restricting the action space of the human hand.
+Although increasing the scale factor can amplify end-effector movement and expand the action space of the robot, it aggravates the difficulty for operators in performing fine-grained tasks.
+These approaches also struggle to handle scenarios where fingers overlap.
+The extra keypoint retargeting module can be computationally expensive in real-time.
 Although specialized hardware [4] can improve accuracy, it substantially increases system costs.
+Our proposed system directly uses the joint pose obtained from the motion capture suit and data gloves, preventing the deviation of vision-based systems and reducing the computational cost.
 Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs.
-Many systems [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control.
-Furthermore, their customization options typically require time-consuming manual parameter tuning that demands significant expertise.
-Figure 3: Pipeline of Our Assisted Module: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution.
-We achieve precise alignment between virtual and physical robots through AprilTag calibration.
+Some approaches [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control.
+Furthermore, your customization options typically require time-consuming manual parameter tuning that requires significant expertise.
 II-B Visual Feedback in Teleoperation Systems
@@
 Traditional approaches primarily rely on real-time camera feeds [24, 25], which can be limited by occlusions and restricted viewing angles.
-Moreover, the lack of preview capabilities for intended commands leaves users (especially novice users) uncertain about the robot’s response, which can increase error rates and user fatigue.
-To address these limitations, researchers have explored enhanced visualization techniques [22, 23, 24] that overlay virtual information onto the workspace view.
-While these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily focus on visualizing task-specific elements.
-This focus on task-specific visualization leaves users uncertain about how the robot will execute its commands.
-III Modular Teleoperation Framework
-Our teleoperation system consists of three main components: wrist pose estimation via IMU-based tracking, hand pose estimation using a motion capture glove, and non-collision retargeting for safe robot command generation.
-The following subsections detail each component of this pipeline.
-III-A Wrist Pose Estimation
-III-A1 SMPL-X Standard
-We adopt the SMPL-X standard [25] as our unified human representation.
+Moreover, the lack of preview capabilities for intended commands leaves users (especially new users) uncertain about the robot’s response, which can increase error rates and user fatigue.
+To address these limitations, researchers have explored enhanced visualization techniques [26, 27, 28, 29, 30] that overlay virtual information on the workspace view.
+Even though these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily design and visualize AR objects for a specific task, such as a pink box with dimensions of 10 cm × 10 cm × 10 cm.
+Any changes to the object’s attributes (e.g., dimensions, color, or type) invalidate the preconfigured AR object, so the system must rebuild it to match the new characteristics.
+Instead of visualizing the manipulated objects, our system provides an AR representation of the robot that directly maps operator commands.
+This design ensures safety for new users who may be unfamiliar with how the robot responds to their inputs.
+Our configuration allows users to use their own URDF file, eliminating the need to reconstruct manipulated objects for every new task.
+III System Overview
+Our system offers an integrated workflow for controlling robot movements with both precision and safety, leveraging low-cost hardware (under \$1,000) and diverse input devices.
+This workflow comprises two complementary modules:
+  1. 1.
+    Teleoperation Module: Includes a teleoperation pipeline for robust motion retargeting.
+  2. 2.
+    Preview-Based Module: Allows operators to check intended movements in a virtual setting before execution.
+As illustrated in Fig. 3, the teleoperation module (Section V) offers a safe and intuitive pipeline that proceeds in three stages:
+  1. 1.
+    Wrist pose estimation via IMU-based tracking (Section IV-A),
+  2. 2.
+    Hand pose estimation using a motion capture glove (Section IV-B),
+  3. 3.
+    Non-collision retargeting to ensure safe joint configurations (Section IV-C).
+Together, these components collectively enable an efficient mapping from human motion to robot commands.
+Moreover, by adopting SMPL-X (Section IV-A1) as a standardized representation of the human body, our pipeline can seamlessly integrate data from any input device conforming to this format.
+Additionally, as illustrated in Fig. 3, the preview-based module (Section V) then provides a safe and intuitive pipeline that proceeds in three stages:
+  1. 1.
+    The preview robot is aligned with the physical robot using AprilTag markers [31] (Section V-B1).
+  2. 2.
+    User commands are visualized on the aligned preview robot (Section V-B2).
+  3. 3.
+    Once the I/O state transitions from active to inactive, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
+This preview-then-execute scheme reduces errors and protects the robot and its surroundings by allowing users to visualize and refine their actions beforehand.
+Thus, it substantially improves efficiency and safety.
+IV Teleoperation Pipeline
+IV-A Wrist Pose Estimation
+Our system tracks the human operator’s wrist motion to enable intuitive robot teleoperation.
+The input consists of raw IMU sensor data from sensors attached to the operator’s arm, while the output is the 6-DoF wrist pose (position and orientation) in world coordinates.
+This wrist pose serves as the primary control signal for the robot’s end-effector, allowing natural mapping between human arm movements and robot motion.
+IV-A1 SMPL-X Standard
+We adopt the SMPL-X standard [32] as our standard representation of the human body.
 The SMPL-X model represents a kinematic tree with standardized joint coordinate systems, where each joint’s pose describes a rotation and translation relative to its parent frame.
-This hierarchical structure enables consistent pose representation across different motion capture devices and simplifies the integration with our TelePhantom system.
-III-A2 IMU-based Tracking System
-While many researchers adopt vision-based motion capture systems in teleoperation, these systems require multiple cameras, precise calibration, and controlled lighting conditions, which limits their practical deployment.
+This hierarchical structure enables consistent pose representation across different motion capture devices and simplifies the integration with our TelePreview system.
+IV-A2 IMU-based Tracking System
 IMU-based motion capture offers robust tracking regardless of visual conditions and occlusions.
-Our system uses IMU sensors located on the user’s palm, forearm, and upper arm.
-Given the three IMUs, we utilize the REBOCAP framework [26] to obtain the specific SMPL-X model of the human body [25, 27].
 Following the SMPL-X convention, we define the world frame system at the midpoint between the feet.
@@
 where $\mathbf{p}_{e}(t)$ and $\mathbf{R}_{e}(t)$ represent the end-effector position and orientation at time $t$, and $\mathbf{p}_{w}(t)$ and $\mathbf{R}_{w}(t)$ represent the user’s wrist position and orientation at time $t$.
-III-B Hand Pose Estimation
-III-B1 Human Hand Joint Value Acquisition
-Vision-based methods typically use deep learning models to predict 3D joint positions $\{p_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ from RGB or RGB-D images.
-Due to the projective nature of imaging, vision-based systems struggle with self-occlusion, forcing researchers to infer joint positions from priors rather than direct observation.
-In contrast, mocap gloves [28] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors despite their higher hardware cost.
-For our teleoperation scenario, we opt for the mocap-based approach to ensure robust hand tracking.
-III-B2 Joint-to-Joint Mapping
+IV-B Hand Pose Estimation
+Mocap gloves [33] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors.
+Our joint-to-joint mapping approach provides a generalizable solution for retargeting high-dimensional human hand motion data to lower-dimensional robotic systems.
+This approach is particularly effective when the input space (e.g., the human hand with 27 DoF) has higher dimensionality than the target space (e.g., the robotic hand with 16 DoF), as it allows selective mapping of the most intuitive and task-relevant degrees of freedom.
+Through manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
 We propose a direct joint-to-joint mapping function $\mathcal{M}:\mathbb{R}^{27}\rightarrow\mathbb{R}^{16}$ that transforms mocap glove readings to LEAP hand [34] joint configurations.
@@
 Where $q_{g}$ and $q_{r}$ represent the joint angles of the glove and the Leap Hand, respectively, and max and min denote their upper and lower joint limits.
-III-C Non-Collision Retargeting Network
-In robotic manipulation systems, ensuring non-collision during operation is crucial for both physical simulator accuracy and hardware safety.
-Self-collisions can lead to unstable behavior in simulators and, more critically, cause hardware damage in real-world deployments.
-This challenge is particularly significant for high-DoF systems like the LEAP hand [29], where collisions often occur due to differences in action spaces and imprecise retargeting.
-As shown in Fig.4, direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig.4 (a).
-While traditional collision avoidance methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
-To address these challenges, we develop a learning-based framework with two key components: (1) a Self-Collision Prediction Network (CPN) that predicts collision probabilities for each joint configuration and (2) a Configuration Correction Network (CCN) that maps invalid configurations to their closest valid counterparts.
-The CCN is trained with a composite loss function that balances configuration similarity and collision avoidance, ensuring both motion continuity and physical feasibility.
-As demonstrated in Fig.
-4 (b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
-We provide the detailed technical implementation of this Pos-to-Pos Non-Collision Module in the Appendix.
-(a) Current retargeting method [6] causing collision.
+(a) Current Retargeting Method [6] Causes Collision.
 (b) Our Self-collision Avoidance Retargeting.
 Figure 4: Comparison of Hand Configuration Retargeting Methods: (a) shows the direct mapping between human and robot hands leading to self-collision; (b) demonstrates our collision-aware retargeting approach that maintains safe configurations.
-IV Phantom System
-The phantom system aims to provide safe and intuitive robot teleoperation through a phantom-based motion preview mechanism.
-The process consists of three key steps: (1) the user’s motion commands are first rendered on a phantom robot, (2) the phantom is precisely aligned with the real robot using AprilTag markers [30], and (3) upon user confirmation via a foot pedal, the validated motion is executed by the physical robot.
-This preview-then-execute workflow, as shown in Fig.
-3, enables users to verify and adjust planned motions before actual execution, significantly improving teleoperation safety and efficiency.
-IV-A Why Phantom System Assistance?
-IV-A1 Current Challenges
+IV-C Non-Collision Retargeting Network
+Direct mapping from human hand postures to robotic hands using existing retargeting approaches like [6] can result in self-collisions, as demonstrated in Fig.
+4(a).
+While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
+To address these challenges, we develop a learning-based framework with two key components:
+  1. 1.
+    A Self-Collision Prediction Network (CPN) that takes robot joint configurations $q\in\mathbb{R}^{n}$ as input and outputs a binary collision probability vector $p\in\mathbb{R}^{m}$ indicating collision likelihood for each link.
+  2. 2.
+    A Configuration Correction Network (CCN) that maps collision-prone joint configurations to their nearest collision-free counterparts.
+Specifically, it transforms an invalid robot configuration $q_{invalid}\in\mathbb{R}^{n}$ to a valid configuration $q_{valid}\in\mathbb{R}^{n}$ that minimizes both collision risk and deviation from the original pose.
+As demonstrated in Fig. 4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
+V Preview Pipeline
+V-A Why Preview System Assistance?
 Traditional teleoperation approaches face significant challenges in data quality when collecting demonstrations for imitation learning.
 During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories.
-In our hanging insertion experiments, we observed that users using direct teleoperation required an average of 4 corrective movements per attempt.
-These adjustments typically involved probing movements to locate the socket edge and multiple approach angle refinements before successful insertion.
-Such trial-and-error behavior introduced significant noise into the collected datasets.
-This issue is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
-IV-A2 Proposed Solution
-Our phantom-assisted teleoperation system addresses these data quality challenges through a novel preview mechanism.
-By visualizing a virtual “phantom” representation of the robot’s intended motion before executing physical movements, users can preview and refine their commands before committing to actual robot motion.
-This preview capability eliminates the need for exploratory movements in the physical space, as the real robot remains stationary until the user confirms the phantom pose.
-This design ensures that recorded trajectories contain only intentional, task-relevant actions without the noise of exploratory movements, directly benefiting downstream imitation learning applications with higher-quality demonstration data.
-IV-B Technical Implementation of Phantom System
-IV-B1 Sim-to-Real Alignment Using AprilTag
-To maintain precise alignment between the phantom and physical robot across multiple camera viewpoints, we implement a robust calibration system using AprilTags [30].
-Our approach establishes reliable reference frames through hand-eye calibration, and AprilTag pose estimation, linking both fixed and floating camera coordinate systems to the robot’s base frame.
+These repeated probing actions to locate the socket edge and refine the approach angle before successful insertion introduced significant noise into the collected datasets.
+Such noise is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
+To address these limitations and provide a more efficient workflow, we propose incorporating a dedicated preview feature that allows users to refine their actions before physical execution.
+By separating exploration from the final command, our approach reduces sub-optimal motions that would otherwise degrade data quality.
+V-B Technical Implementation of Preview System
+We design our preview feature to integrate seamlessly with the teleoperation workflow.
+It requires three critical elements: (1) precise spatial calibration to maintain alignment between the virtual preview and the physical robot, (2) realistic rendering to blend the preview with camera feeds, and (3) state management to handle transitions between preview and execution phases.
+The following subsections detail each of these technical components, highlighting how they collectively enable users to view and adjust planned motions prior to committing them to the physical robot.
+V-B1 Spatial Alignment Using AprilTag
+To maintain a precise spatial alignment between the preview and the physical robot in the camera viewpoints, we implement a robust calibration system using AprilTags [31].
+Our approach involves two sequential steps: (1) a hand-eye calibration that establishes the transformation between the robot’s base frame and the fixed camera frame, and (2) AprilTag pose detection that enables real-time tracking of the robot’s position and orientation relative to each camera view.
+This chain of transformations ensures consistent preview visualization across different viewpoints.
 Figure 5: Our Transformation Relationship: The number in the circle denotes the order of transformation acquisition.
@@
 For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3).
-This chain of transformations ensures consistent phantom visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
-IV-B2 Robot Phantom Visualization
-Building upon this spatial alignment system, we implement the visual component of our phantom system.
+This chain of transformations ensures consistent preview visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
+V-B2 Visual Preview Visualization
+Building upon this spatial alignment system, we implement the visual component of our preview system.
 As shown in Fig.
 6(a), we use the pyrender [35] library to render the virtual robot through the previous float camera.
-The rendered phantom image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment.
-This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the phantom accurately reflects the robot’s intended configuration in the workspace.
+The rendered preview image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment.
+This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the preview accurately reflects the robot’s intended configuration in the workspace.
 The resulting overlay provides users with an intuitive preview of planned motions.
-IV-B3 Multi-view Phantom Visualization
-Task TelePhantom Open Teach [3] AnyTeleop[6] Telekinesis[32]
+Our preview system also supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks.
+As shown in Fig.
+6(b) and Fig.
+6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives.
+Task TelePreview Open Teach [3] AnyTeleop[6] Telekinesis[36]
 Pick & Place 1.0 0.8 1.0 0.9
@@
 TABLE I: Real Robot Teleoperation Results.
-For the baseline methods, we use success rate as a metric reported in their papers.
-Especially, we use the success rate of an expert as a metric for Open Teach [3].
-(a) Real-time phantom rendering
+We use the success rates the baseline methods report in their papers, marking tasks they did not attempt or do not support with “-”.
+Success rates for Open Teach [3] reflect expert performance.
+(a) Preview Rendering
 (b) Third-person View
@@
 Task Success Rate Average Success Execution Time (s)
-w/o phantom w/ phantom Difference $\uparrow$ w/o phantom w/ phantom Difference $\downarrow$
+w/o preview w/ preview Difference $\uparrow$ w/o preview w/ preview Difference $\downarrow$
 Pick & Place 0.6 1.0 +0.4 23.56 13.55 -10.01
@@
 Cup Stacking 0.5 1.0 +0.5 31.54 18.91 -12.63
-TABLE II: Effect of Phantom Assistance on New User Performance.
-Our phantom system supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks.
-As shown in Fig.
-6(b) and Fig.
-6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives.
-For instance, during manipulation tasks, users can simultaneously monitor the robot’s vertical position from the third-person view while checking its planar alignment from the top-down view.
-IV-B4 IO Control for Phantom Visualization
-Figure 7: State transition of phantom control
+TABLE II: Effect of Preview Assistance on New User Performance.
+V-B3 IO Control for Preview Visualization
 Our system uses a foot pedal as an intuitive control interface for state transitions.
 As shown in Fig.
-7, when the IO is activated, the physical robot stops while the phantom appears and responds to user commands for motion preview.
-Upon IO deactivation, the phantom disappears, and the system extracts its final configuration as the target pose, bypassing any exploratory movements made during the preview.
+7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode).
+Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode).
 This target is then processed by the mplib motion planning library [37] to generate an optimal trajectory for execution.
@@
 This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
-V Experiments
+Figure 7: State Transition of Preview Control
+VI Experiments
 Our experiments aim to address the following questions:
-(Q1) How effective is TelePhantom vs Baselines?
-(Q2) How beneficial is TelePhantom for new users?
-(Q3) How versatile is TelePhantom on robots?
-V-A Experiment Setup
-TelePhantom integrates several key components.
-The user’s movements are captured via an inertial motion capture suit and a pair of strain gauge-based mocap gloves, providing high-fidelity joint angle data for both body posture and delicate hand movements.
-To visualize robot action before execution, we employ a RealSense RGB-D camera and render the phantom either on a standard 2D display or within a VR headset (e.g., Meta Quest 3).
-For the physical hardware, we use a UFactory xArm-6 robotic arm outfitted with a Leap hand [29], utilizing AprilTag [30] to align the initial poses of the phantom and the real robot.
+Q1: How effective is TelePreview vs. Baselines?
+Q2: How beneficial is TelePreview for new users?
+Q3: How adaptable is TelePreview for new hardware?
+VI-A Experiment Setup
+TelePreview integrates several key components.
+We capture the user’s movements via an inertial motion capture suit [38] and a pair of strain-gauge-based mocap gloves [33], providing high-fidelity joint angle data for both body posture and delicate hand movements.
+To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3[39]).
+For the physical hardware, we use a UFactory xArm-6 robotic arm[40] outfitted with a Leap hand [34], utilizing AprilTag [31] to align the initial poses of the preview and the real robot.
 Following the methodology outlined in [36, 6], users attempted each task 10 times. We reference the baseline success rates from their papers.
-V-B Real Robot Teleoperation
-We test TelePhantom on five real-world tasks that highlight different manipulation challenges:
+VI-B Task Descriptions
+We test TelePreview on five real-world tasks, each highlighting different manipulation challenges:
   • Pick & Place: Grasp an object and place it at precise locations.
@@
   • Cup Stacking: Stack a cup onto another cup.
-Performance Comparison with Baselines.
-Addressing Q1 (effectiveness vs.
-baselines), we compare the real robot teleoperation performance of TelePhantom, operated by an expert, with Telekinesis [32], AnyTeleop [6] and Open Teach [3].
-As summarized in Table I, TelePhantom outperforms baseline methods in terms of success rate.
-The baseline methods are limited to simple, short-horizon tasks.
-While the baseline methods can handle basic tasks, they fail to accurately capture subtle wrist rotations and movements in more delicate operations like the hung task.
-The superior performance of TelePhantom in these scenarios demonstrates its capability to handle more challenging tasks effectively.
-Effect of Phantom.
-To address Q2 (benefits for new users), we assess TelePhantom’s utility by measuring both the success rate and the average execution time on the real robot (demonstration time in imitation learning) with and without phantom.
-As shown in Table II, TelePhantom significantly improves the performance of new users by improving their task success rate.
-The execution time on a real robot corresponds to the demonstration time required for imitation learning.
-High-quality demonstrations are typically associated with shorter execution times.
-With the assistance of TelePhantom, new users achieve a lower average execution time, indicating that they can produce higher-quality demonstrations compared to users without access to TelePhantom.
-For new users, command visualization and reduced execution time not only decrease physical strain but also improve safety.
-V-C Deployment on Various Robots
-TelePhantom’s modular design enables seamless adaptation to different input devices and end-effectors.
-Thanks to our robot-centric phantom approach and standardized transformation framework, users can readily configure the system for various robotic platforms by simply updating the kinematic parameters and transformation chain.
-This adaptability across different robot configurations, as posed in Q3, demonstrates TelePhantom’s versatility as a flexible teleoperation solution that can serve diverse robotic applications with minimal modification requirements.
-VI Conclusion
-In this paper, we presented TelePhantom, a low-cost teleoperation system featuring AR phantom feedback designed for deployment across diverse robotic platforms.
-Our experiments demonstrate that TelePhantom offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks.
-Through its innovative phantom visualization approach, the system provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
-While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the phantom and scene objects.
+VI-C Evaluation Metrics
+We measure two key metrics for each task:
+  • Success Rate: We define the success rate as the proportion of times the operator completes the task within the given constraints.
+For example, if the user attempts a cup-hanging task $X$ times and successfully hangs the cup on the rack $Y$ times, the success rate is $\frac{Y}{X}$.
+  • Execution Time: The total duration from task initiation to successful completion (i.e., the length of the recorded demo episode).
+VI-D Performance Comparison with Baselines (Q1)
+To address Q1—evaluating how our approach compares to existing teleoperation methods—we measure the actual robot performance of TelePreview against three baselines: Telekinesis [36], AnyTeleop [6], and Open Teach [3].
+For fairness and consistency, an expert user operates each method and tests it on the five tasks described in Section VI-B.
+The primary metrics are success rate and average execution time.
+As summarized in Table I, TelePreview achieves higher success rates than all baselines.
+While Telekinesis, AnyTeleop, and Open Teach handle simpler, short-horizon tasks, they often struggle with complex manipulations requiring subtle wrist rotations and fine-grained positioning (e.g., the Hang task).
+TelePreview, by contrast, captures these nuances more effectively, reflecting stronger overall performance and versatility.
+These findings demonstrate TelePreview’s ability to address a broad spectrum of real-world challenges, from basic pick-and-place operations to more intricate tasks involving delicate movements.
+(a) LEAP Hand
+(b) Gripper
+Figure 8: Deployment on Different Robots.
+VI-E Effect of Preview for Users (Q2)
+To investigate Q2—how the preview feature benefits new (non-expert) users—we evaluate both success rate and average execution time in two conditions: (1) with the preview enabled and (2) with the preview disabled.
+The same five tasks from Section VI-B are attempted by new users, with each task repeated 10 times.
+As shown in Table II, enabling the preview boosts user performance.
+Success rates increase across all tasks, with especially notable gains in challenging manipulations such as cup stacking (+0.5) and pick & place (+0.4).
+Moreover, average execution times drop by up to 12.63 seconds (cup stacking), demonstrating not only improved accuracy but also greater efficiency.
+By visualizing a virtual “preview” of the robot’s intended motion and confirming the final pose prior to executing physical movements, users eliminate the need for exploratory motions in the actual workspace.
+This design ensures that recorded trajectories contain only intentional, task-oriented actions without extraneous trial-and-error, providing cleaner, high-quality demonstrations for imitation learning.
+As a result, TelePreview delivers both a more intuitive teleoperation experience and more consistent data for downstream training.
+VI-F Deployment on Various Robots (Q3)
+We deploy our TelePreview system on a Ufactory xArm [40] equipped with both a LEAP hand [34] and a parallel gripper.
+By updating only the kinematic parameters and transformation chain for each end-effector, TelePreview preserves its full functionality across diverse setups, as shown in Fig.
+8.
+This successful deployment on multiple end-effector configurations illustrates the system’s adaptability to new hardware.
+Its modular design allows users to seamlessly integrate TelePreview into different robotic platforms while retaining all key features—preview, multi-view visualization, and intuitive control interfaces.
+VII Conclusion
+In this paper, we presented TelePreview, a low-cost teleoperation system featuring AR preview feedback designed for deployment across diverse robotic platforms.
+Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks.
+TelePreview provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
+While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the preview robot and scene objects.
 Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.
-Addressing these visualization challenges could yield several essential benefits.
-Improved occlusion handling would reduce user uncertainty during complex manipulation tasks, leading to higher-quality demonstration data for imitation learning.
-Additionally, more accurate spatial rendering would create a more intuitive teleoperation experience, potentially reducing user training time and cognitive load.
-These improvements would further advance TelePhantom’s goal of providing accessible, effective teleoperation capabilities across diverse robotic applications.

### v6

@@
 We demonstrate that it outperforms other teleoperation systems across five tasks, emphasize its ease of use, and highlight its straightforward deployment across diverse robotic platforms.
-We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview/.
+We release our code and a deployment document on our website https://nus-lins-lab.github.io/telepreview-web/
 {strip}
@@
 I Introduction
-Figure 2: Overview of the TelePreview System Architecture: Our system consists of three main components: (1) Various input devices, including RGB-D cameras, IMU mocap suits, VR headsets, mocap gloves, AR hand tracking, and EMG armbands for capturing human motion; (2) A central processing pipeline that uses SMPL-X as a standard to process wrist and hand pose data, feeding into our Preview System with joint-to-joint mapping and a non-collision retargeting network; (3) Support for different robot platforms as output devices for executing the mapped movements.
-Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+Figure 2: Overview of the System Architecture (1) Various input devices for capturing human motion.
+Only one device from each of the two input groups (wrist pose and hand gesture) is required during operation.
+The VR headset is shown at the intersection of the two groups, as it can capture both; (2) A processing pipeline based on SMPL-X that performs joint mapping and collision-free retargeting; (3) Output to various robot platforms for executing the mapped motions.
+Researchers have long recognized teleoperation as an essential component for gathering on-robot data used in learning from demonstrations [1, 2, 3, 4, 5, 6].
 An intuitive and user-friendly teleoperation system is crucial for collecting high-quality, diverse, and scalable data.
 However, effectively controlling robots with dexterous hands remains a significant challenge.
-The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex.
+The fundamental differences between human and robotic hands make the direct mapping of human movements to robotic actions exceptionally complex [7, 8, 9].
 These inherent structural and functional disparities, combined with the precision required for fine-grained tasks, often result in inefficient and potentially compromised data collection processes.
@@
 II Related Work
-Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides visual preview before physical execution.
+Figure 3: Pipeline of Our Modules: The system tracks user wrist and hand poses, maps them to robot joint configurations through joint-to-joint mapping and non-collision retargeting, and provides a visual preview before physical execution.
 We achieve precise alignment between virtual and physical robots through AprilTag calibration.
-II-A Robot Teleoperation Frameworks for Manipulation
-The growing interest in training robots through imitation learning has driven the need for extensive, high-quality real robot datasets.
+II-A Teleoperation Frameworks for Dexterous Manipulation
 Teleoperation provides an efficient method for demonstrating and recording intricate robotic tasks[10, 11, 12].
-Researchers have explored various teleoperation systems, including VR controllers [1, 2, 3], motion capture [16, 17, 18], wearable gloves [19, 11, 20, 21], exoskeletons [22, 23], each offering distinct benefits in terms of accessibility, precision, and generalizability.
-Vision-based teleoperation systems [6] typically employ RGB or RGB-D cameras to detect hand poses.
-The hand pose retargeting module maps the human hand pose data obtained from perception algorithms into joint positions of the dexterous hand.
-We formulate this process as an optimization problem that minimizes the difference between the key point vectors of the human hand and the dexterous hand.
-To generate the end-effector motions, 3D positions of key points in the camera frame are calculated by using an RGB-D camera.
-The wrist pose is then computed by keypoint positions in the local wrist frame and global camera frame using the Perspective-n-Point (PnP) algorithm.
-Although vision-based teleoperation systems offer a relatively low-cost solution [6, 2, 3, 1], they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspace, and high computational demands.
-This limitation stems from the requirement of vision-based approaches for operators to keep their hands within the camera’s field of view, thereby restricting the action space of the human hand.
-Although increasing the scale factor can amplify end-effector movement and expand the action space of the robot, it aggravates the difficulty for operators in performing fine-grained tasks.
-These approaches also struggle to handle scenarios where fingers overlap.
-The extra keypoint retargeting module can be computationally expensive in real-time.
+But Many existing teleoperation frameworks focus on low-degree-of-freedom (low-DoF) robotic grippers, using input modalities such as keyboards[13], joysticks[14, 15], or gamepads[16] to issue discrete commands.
+While these are effective for simple tasks, they are not sufficient for the fine-grained control of dexterous robotic hands with High DoFs.
+Researchers have explored various teleoperation systems for dexterous manipulation, including VR headsets [7, 17, 1], motion capture [18, 19, 20], wearable gloves [21, 6, 22, 23, 24], exoskeletons [25, 26].
+Using VR headsets, vision-based teleoperation systems [3, 17, 1, 7] typically employ RGB cameras to detect hand poses.
+These systems often rely on estimated key points from visual perception to map user motions onto the robot, typically using optimization-based retargeting methods.
+Although vision-based teleoperation systems offer a relatively low-cost solution, they face significant challenges, including limited degrees of freedom (DoFs), constrained reachable workspaces, and high computational demands.
+These systems also struggle with occlusions, finger overlap, and other perception issues.
+The keypoint-based retargeting step can be computationally expensive in real time.
 Although specialized hardware [2] can improve accuracy, it substantially increases system costs.
-Our proposed system directly uses the joint pose obtained from the motion capture suit and data gloves, preventing the deviation of vision-based systems and reducing the computational cost.
-Recent unified teleoperation frameworks attempt to address these limitations but often introduce new trade-offs.
-Some approaches [3, 6] prioritize implementation simplicity by tracking only fingertip positions rather than enabling detailed joint-level control.
-Furthermore, your customization options typically require time-consuming manual parameter tuning that requires significant expertise.
+In contrast, exoskeleton systems[25, 26] provide highly accurate motion capture with low latency but tend to be tightly coupled to specific hardware platforms, which limits their generalizability across different robot embodiments.
+In this work, we adopt a glove-based teleoperation framework that captures human hand motion at the joint level and maps it to a unified body model (e.g., SMPL-X) before retargeting it to the robot.
+This abstraction layer enables real-time, joint-level control of dexterous robotic hands while improving generalizability across different robot embodiments.
+Although glove-based teleoperation has previously been considered expensive, recent commercialization and the availability of low-cost, off-the-shelf data gloves have made this approach increasingly accessible.
 II-B Visual Feedback in Teleoperation Systems
@@
 Moreover, the lack of preview capabilities for intended commands leaves users (especially new users) uncertain about the robot’s response, which can increase error rates and user fatigue.
-To address these limitations, researchers have explored enhanced visualization techniques [26, 27, 28, 29, 30] that overlay virtual information on the workspace view.
-Even though these approaches improve user perception by displaying planned trajectories and environmental constraints, they primarily design and visualize AR objects for a specific task, such as a pink box with dimensions of 10 cm × 10 cm × 10 cm.
-Any changes to the object’s attributes (e.g., dimensions, color, or type) invalidate the preconfigured AR object, so the system must rebuild it to match the new characteristics.
-Instead of visualizing the manipulated objects, our system provides an AR representation of the robot that directly maps operator commands.
-This design ensures safety for new users who may be unfamiliar with how the robot responds to their inputs.
-Our configuration allows users to use their own URDF file, eliminating the need to reconstruct manipulated objects for every new task.
+To address these limitations, researchers have explored enhanced visualization techniques [29, 30, 31, 32] in AR-based teleoperation systems[33, 34], which aim to improve user understanding of spatial constraints and planned motions.
+Prior AR-based teleoperation systems typically display a virtual robot arm in the environment but do not spatially align it with the physical robot.
+As a result, they cannot capture interaction data during contact with real objects or the robot’s behavior following physical execution.
+In contrast, our system overlays the AR robot directly on top of the physical robot, enabling users to first control a virtual preview and then switch to executing the same motion with the real hardware.
+This design allows us to capture the full sequence of a teleoperation task—from intent to contact to post-contact behavior—without requiring real-to-sim translation or object re-modeling.
+By eliminating the need to reconstruct physical objects in simulation, we reduce system setup complexity while enabling high-fidelity, real-world demonstration collection.
 III System Overview
@@
   3. 3.
-    Once the I/O state transitions from active to inactive, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
+    Once the I/O state transitions, the final preview pose is captured and converted into an optimal trajectory for actual execution (Section V-B3).
 This preview-then-execute scheme reduces errors and protects the robot and its surroundings by allowing users to visualize and refine their actions beforehand.
@@
 IV-B Hand Pose Estimation
-Mocap gloves [33] offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors.
+Mocap gloves offer direct joint angle measurements $\{q_{i}\}_{i=1}^{27}$ using flex sensors.
 Our joint-to-joint mapping approach provides a generalizable solution for retargeting high-dimensional human hand motion data to lower-dimensional robotic systems.
 This approach is particularly effective when the input space (e.g., the human hand with 27 DoF) has higher dimensionality than the target space (e.g., the robotic hand with 16 DoF), as it allows selective mapping of the most intuitive and task-relevant degrees of freedom.
-Through manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
+Through a manual selection of corresponding joints based on human intuition and task requirements, followed by range alignment between input and output spaces, we maintain natural and predictable motion correspondence.
 We propose a direct joint-to-joint mapping function $\mathcal{M}:\mathbb{R}^{27}\rightarrow\mathbb{R}^{16}$ that transforms mocap glove readings to LEAP hand [37] joint configurations.
@@
 4(a).
-While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations.
+While traditional retargeting methods can achieve real-time performance for simpler end-effectors, they rely on optimization-based approaches that become computationally intensive for high-DoF configurations, highlighting the importance of collision-aware methods such as [32].
 To address these challenges, we develop a learning-based framework with two key components:
@@
 Specifically, it transforms an invalid robot configuration $q_{invalid}\in\mathbb{R}^{n}$ to a valid configuration $q_{valid}\in\mathbb{R}^{n}$ that minimizes both collision risk and deviation from the original pose.
-As demonstrated in Fig. 4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
+As demonstrated in Fig.
+4(b), our method successfully transforms problematic configurations into stable, safe positions while maintaining efficient real-time performance (about 60Hz).
+The implementation detail can be found in the Appendix.
 V Preview Pipeline
@@
 During complex manipulation tasks, users frequently make exploratory movements that contaminate the demonstration data with sub-optimal trajectories.
-These repeated probing actions to locate the socket edge and refine the approach angle before successful insertion introduced significant noise into the collected datasets.
 Such noise is particularly problematic for training imitation learning models, which benefit most from clean, purposeful demonstrations rather than trajectories cluttered with exploratory movements.
+Recent research has highlighted how errors or exploratory movements in teleoperation can decrease the generalization ability of the learned policies and negatively impact performance on unseen tasks.[38, 39, 40, 41, 42]
 To address these limitations and provide a more efficient workflow, we propose incorporating a dedicated preview feature that allows users to refine their actions before physical execution.
@@
 V-B Technical Implementation of Preview System
-We design our preview feature to integrate seamlessly with the teleoperation workflow.
+We designed our preview feature to integrate seamlessly with the teleoperation workflow.
 It requires three critical elements: (1) precise spatial calibration to maintain alignment between the virtual preview and the physical robot, (2) realistic rendering to blend the preview with camera feeds, and (3) state management to handle transitions between preview and execution phases.
@@
 Our approach involves two sequential steps: (1) a hand-eye calibration that establishes the transformation between the robot’s base frame and the fixed camera frame, and (2) AprilTag pose detection that enables real-time tracking of the robot’s position and orientation relative to each camera view.
-This chain of transformations ensures consistent preview visualization across different viewpoints.
 Figure 5: Our Transformation Relationship: The number in the circle denotes the order of transformation acquisition.
@@
 5, our calibration procedure involves multiple sequential transformations.
-We begin with hand-eye calibration of the fixed top-view camera (1), followed by detecting the AprilTag’s pose from this camera’s perspective (2).
-Using these relationships, we can compute the transformation between the AprilTag and the robot base frame (2).
-For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3).
-This chain of transformations ensures consistent preview visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
+We begin with hand-eye calibration of the fixed top-view camera, followed by AprilTag pose detection from this camera’s perspective.
+These steps allow us to compute the transformation from the AprilTag to the robot base frame.
+For the floating third-view camera, we detect the AprilTag pose and use it to determine its position relative to the robot base frame.
+This calibration ensures consistent visualization across dynamic or multi-view setups.
 V-B2 Visual Preview Visualization
@@
 As shown in Fig.
-6(a), we use the pyrender [35] library to render the virtual robot through the previous float camera.
-The rendered preview image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment.
-This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the preview accurately reflects the robot’s intended configuration in the workspace.
-The resulting overlay provides users with an intuitive preview of planned motions.
-Our preview system also supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks.
+6(a), we use the pyrender library to render the virtual robot through the float camera.
+The preview image is then composited with the real camera feed using alpha blending, creating an overlay where the virtual robot appears aligned with the physical environment.
+This requires the calibration of virtual camera parameters to match the physical cameras, ensuring the preview accurately reflects the robot’s intended configuration.
+The preview offers real-time visual guidance before execution.
+Our system supports multiple viewpoints to aid spatial reasoning and trajectory checking.
 As shown in Fig.
 6(b) and Fig.
-6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives.
+6(c), we deploy third-person and top-down cameras, rendered and shown simultaneously on an external monitor.
+The system does not require a VR headset.
+For instance, during manipulation tasks, users can simultaneously monitor the robot’s vertical position from the third-person view while checking its planar alignment from the top-down view.
+V-B3 IO Control for Preview Visualization
+Our system uses a foot pedal as an intuitive control interface for state transitions.
+As shown in Fig.
+7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode).
+Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode).
+The mplib motion planning library then processes this target to generate an optimal trajectory for execution.
+During this motion, the user does not need to remain static—the robot executes the planned motion autonomously.
+Once the target pose is reached, the system enters a continuous teleoperation mode in which the robot follows the user’s live movements in real time until the next preview is initiated.
+This workflow ensures that users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
+(a) Preview Rendering
+(b) Third-person View
+(c) Top-down View
+Figure 6: Rendering and Multi-view Visualization System.
+Figure 7: State Transition of Preview Control
 Task TelePreview Open Teach [1] AnyTeleop[3] Telekinesis[32]
@@
 Cup Stacking 1.0 - 0.7 0.3
-TABLE I: Real Robot Teleoperation Results.
-We use the success rates the baseline methods report in their papers, marking tasks they did not attempt or do not support with “-”.
-Success rates for Open Teach [3] reflect expert performance.
-(a) Preview Rendering
-(b) Third-person View
-(c) Top-down View
-Figure 6: Rendering and Multi-view Visualization System.
+TABLE I: Real Robot Teleoperation Results We mark tasks they did not attempt or do not support with “-”. Success rates for Open Teach [1] reflect expert performance.
 Task Success Rate Average Success Execution Time (s)
 w/o preview w/ preview Difference $\uparrow$ w/o preview w/ preview Difference $\downarrow$
-Pick & Place 0.6 1.0 +0.4 23.56 13.55 -10.01
-Hang 0.6 1.0 +0.4 29.30 30.83 +1.53
-Pour 0.9 0.8 -0.1 43.20 36.13 -7.07
-Box Rotation 0.6 0.8 +0.2 30.47 19.12 -11.35
-Cup Stacking 0.5 1.0 +0.5 31.54 18.91 -12.63
+Pick & Place 0.6 1.0 +0.4 23.56±$\pm 4.65$ 13.55±$\pm 3.17$10.01
+Hang 0.6 1.0 +0.4 29.30±$\pm 8.18$ 30.83±$\pm 3.65$ +1.53
+Pour 0.9 0.8 -0.1 43.20±$\pm 6.73$ 36.13±$\pm 5.86$7.07
+Box Rotation 0.6 0.8 +0.2 30.47±$\pm 14.08$ 19.12±$\pm 3.32$11.35
+Cup Stacking 0.5 1.0 +0.5 31.54±$\pm 6.04$ 18.91±$\pm 1.29$12.63
 TABLE II: Effect of Preview Assistance on New User Performance.
-V-B3 IO Control for Preview Visualization
-Our system uses a foot pedal as an intuitive control interface for state transitions.
-As shown in Fig.
-7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode).
-Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode).
-This target is then processed by the mplib motion planning library [37] to generate an optimal trajectory for execution.
-After reaching the target pose, the robot automatically resumes real-time teleoperation until the next IO is activated.
-This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
-Figure 7: State Transition of Preview Control
 VI Experiments
@@
 TelePreview integrates several key components.
-We capture the user’s movements via an inertial motion capture suit [38] and a pair of strain-gauge-based mocap gloves [33], providing high-fidelity joint angle data for both body posture and delicate hand movements.
-To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3[39]).
-For the physical hardware, we use a UFactory xArm-6 robotic arm[40] outfitted with a Leap hand [34], utilizing AprilTag [31] to align the initial poses of the preview and the real robot.
+We capture the user’s movements via an inertial motion capture suit and a pair of strain-gauge-based mocap gloves, providing high-fidelity joint angle data for both body posture and delicate hand movements.
+To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3).
+For the physical hardware, we use a UFactory xArm-6 robotic arm outfitted with a Leap hand [37], utilizing AprilTag [35] to align the initial poses of the preview and the real robot.
 Following the methodology outlined in [32, 3], users attempted each task 10 times. We reference the baseline success rates from their papers.
 VI-B Task Descriptions
-We test TelePreview on five real-world tasks, each highlighting different manipulation challenges:
+We test TelePreview on five real-world tasks.
+These tasks were carefully selected to cover a diverse range of manipulation skills, including grasping, rotation, alignment, and insertion, and are representative of common benchmarks used in teleoperation research:
   • Pick & Place: Grasp an object and place it at precise locations.
@@
   • Success Rate: We define the success rate as the proportion of times the operator completes the task within the given constraints.
-For example, if the user attempts a cup-hanging task $X$ times and successfully hangs the cup on the rack $Y$ times, the success rate is $\frac{Y}{X}$.
   • Execution Time: The total duration from task initiation to successful completion (i.e., the length of the recorded demo episode).
 VI-D Performance Comparison with Baselines (Q1)
-To address Q1—evaluating how our approach compares to existing teleoperation methods—we measure the actual robot performance of TelePreview against three baselines: Telekinesis [36], AnyTeleop [6], and Open Teach [3].
-For fairness and consistency, an expert user operates each method and tests it on the five tasks described in Section VI-B.
-The primary metrics are success rate and average execution time.
+Due to frequent self-collision failures encountered during local deployment of vision-based methods, we report results from their original publications and discuss reproduction experiences in the Appendix.
 As summarized in Table I, TelePreview achieves higher success rates than all baselines.
@@
 TelePreview, by contrast, captures these nuances more effectively, reflecting stronger overall performance and versatility.
-These findings demonstrate TelePreview’s ability to address a broad spectrum of real-world challenges, from basic pick-and-place operations to more intricate tasks involving delicate movements.
-(a) LEAP Hand
-(b) Gripper
-Figure 8: Deployment on Different Robots.
 VI-E Effect of Preview for Users (Q2)
 To investigate Q2—how the preview feature benefits new (non-expert) users—we evaluate both success rate and average execution time in two conditions: (1) with the preview enabled and (2) with the preview disabled.
+Our study involved 10 participants in total, including 5 new users with no prior experience in robotic teleoperation and 5 experienced users (3 of whom are authors).
+This section focuses on the new user group.
+Participants were given consistent task instructions and evaluated under a standardized protocol, including predefined success criteria.
+Task order was counterbalanced across participants to mitigate ordering effects.
 The same five tasks from Section VI-B are attempted by new users, with each task repeated 10 times.
+We report both mean and standard deviation across trials for each task.
+In addition to objective performance, we collected subjective workload ratings using a 5-point Likert scale (using the same evaluation metrics in [43]) and recorded the practice time required for participants to reach task readiness under different input modalities.
+Detailed results and analysis are provided in the Appendix.
 As shown in Table II, enabling the preview boosts user performance.
@@
 Moreover, average execution times drop by up to 12.63 seconds (cup stacking), demonstrating not only improved accuracy but also greater efficiency.
+The standard deviations across trials are also reduced, suggesting more consistent user behavior.
 By visualizing a virtual “preview” of the robot’s intended motion and confirming the final pose prior to executing physical movements, users eliminate the need for exploratory motions in the actual workspace.
-This design ensures that recorded trajectories contain only intentional, task-oriented actions without extraneous trial-and-error, providing cleaner, high-quality demonstrations for imitation learning.
-As a result, TelePreview delivers both a more intuitive teleoperation experience and more consistent data for downstream training.
+This design ensures that recorded trajectories contain only intentional, task-oriented actions .
+This results in higher-quality demonstrations that are free from unnecessary trial-and-error behaviors.
 VI-F Deployment on Various Robots (Q3)
-We deploy our TelePreview system on a Ufactory xArm [40] equipped with both a LEAP hand [34] and a parallel gripper.
-By updating only the kinematic parameters and transformation chain for each end-effector, TelePreview preserves its full functionality across diverse setups, as shown in Fig.
-8.
-This successful deployment on multiple end-effector configurations illustrates the system’s adaptability to new hardware.
-Its modular design allows users to seamlessly integrate TelePreview into different robotic platforms while retaining all key features—preview, multi-view visualization, and intuitive control interfaces.
+We evaluate the generalizability of TelePreview across three end-effector types on a Ufactory xArm: a 16-DoF anthropomorphic LEAP hand, a parallel-jaw gripper, and a vacuum gripper (both 1-DoF).
+Our results (see Appendix) show that TelePreview yields the most significant performance gains with the LEAP hand, highlighting its effectiveness for high-DoF, fine-grained manipulation tasks.
+Improvements with the gripper and vacuum gripper are more modest but still demonstrate the system’s utility across diverse hardware configurations.
+These findings support our claim that TelePreview generalizes well while providing the greatest benefit in dexterous settings.
 VII Conclusion
 In this paper, we presented TelePreview, a low-cost teleoperation system featuring AR preview feedback designed for deployment across diverse robotic platforms.
-Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks.
+Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to perform a variety of fine-grained tasks successfully.
 TelePreview provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
 While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the preview robot and scene objects.
-Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.
+This limitation stems from our use of alpha blending, which lacks depth-aware occlusion reasoning and cannot distinguish whether the virtual arm should occlude real objects or vice versa, resulting in artifacts even on passthrough-capable devices like the Meta Quest 3 without real-time depth estimation.
+To mitigate this ambiguity, our approach employs multi-view overlays (e.g., top-down and side views), allowing operators to compare perspectives and correctly infer spatial ordering.
+Future work will explore incorporating depth estimation from RGB-D or passthrough feeds and using it to inform depth-aware rendering and spatial reasoning.
+These could reduce user uncertainty during manipulation, leading to cleaner demonstration data and a more intuitive teleoperation experience.